A student measures the time T for one complete oscillation of a pendulum of length l.
Her results are shown below.
l/m-> 0.420+_ 0.001
T/s-> 1.3+_ 0.1
She uses the formula
T=2π√l÷g
to calculate the acceleration of free fall g.
What is the best estimate of tje percentage uncertainty in the value of g?
A)0.02%
B)4%
C)8%
D)16%
1
Expert's answer
2020-11-16T07:52:00-0500

Expressing "g" from the formula, obtain:


"g = \dfrac{4\pi^2l}{T^2}"

Let's differentiate "g":


"dg = \dfrac{4\pi^2}{T^2}dl - \dfrac{8\pi^2l}{T^3}dT"

where "dl = 0.001m" is the uncertainty in length and "dT = 0.1s" is the uncertainty in time. The percentage uncertainty in "g" is then


"\delta=\left|\dfrac{dg}{g}\right| = \left|\dfrac{dl}{l} - 2\dfrac{dT}{T}\right|"

The maximum uncertainty occurs when "dT<0". Then:


"\delta_{max}=\left|\dfrac{dl}{l}\right| + 2\left|\dfrac{dT}{T}\right|\\
\delta_{max}=\left|\dfrac{0.001}{0.420}\right| + 2\left|\dfrac{0.1}{1.3}\right| \approx 0.156 \approx 16\%"

Answer. 16%.